OpenStudy (anonymous):

can some one explain to me what the difference is when they both look the same? how can i tell the difference?

OpenStudy (anonymous):

\[|x+4|>-12 final answer is (-\infty,\infty)\]

OpenStudy (anonymous):

\[|x+1|>-3 which when worked out is (-\infty,-4)\cup(2,\infty)\]

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

would you agree that \[|(x+4)|=|-(x+4)|\]

OpenStudy (anonymous):

well the left side would need to be added unlike the right side

OpenStudy (amistre64):

also, when is the absolute value function ever less than 0?

OpenStudy (anonymous):

it isn't

OpenStudy (amistre64):

the function y = |x+k|, has the same range as y = |x|, its just moved to the left or right, not up or down

OpenStudy (anonymous):

oh i see

OpenStudy (amistre64):

\[|x+4|>-12~:~(-\infty,,\infty)\] \[|x+1|>-3~:~(-\infty,,\infty)\]